# MLLMEraser: Achieving Test-Time Unlearning in Multimodal Large Language Models through Activation Steering

### v2

@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023; Wang et al., 2024; Zhu et al., 2024; Yang et al., 2023; Anil et al., 2023) have demonstrated remarkable capabilities in tasks such as visual question answering (Hu et al., 2024; Kuang et al., 2025), image–text generation (Wu et al., 2024b; Lan et al., 2025), and embodied AI applications (Wu et al., 2024a; Cheng et al., 2025). However, their large-scale deployment raises concerns about memorizing problematic information once learned, particularly in privacy-sensitive (Huo et al., 2025) or safety-critical (Liu et al., 2024c) applications, highlighting the need for reliable unlearning mechanisms to ensure trustworthy MLLM systems (Li et al., 2024; Liu et al., 2025c; Chen et al., 2025). MLLM unlearning aims to selectively erase designated information across modalities while preserving general utility, thereby supporting privacy protection, mitigating misuse, and maintaining reliability (Dontsov et al., 2025). Existing work mainly adapts training–based strategies from LLM unlearning, employing gradient ascent, preference optimization (Zhang et al., 2024), or performing targeted parameter updates (Liu et al., 2025d; Huo et al., 2025). While effective, these training-based methods introduce substantial computational costs, inference latency, and risks of corrupting retained knowledge (Ding et al., 2025). This motivates test-time MLLM unlearning (See Figure 1), a paradigm that prevents the generation of designated information at inference without modifying model parameters, which offers an immediate, lightweight, and reversible solution.
@@ 2 Preliminary
This section begins by introducing the notation and formalizing the problem of test-time MLLM unlearning in Section 2.1. Section 2.2 outlines how activation steering offers a feasible mechanism for test-time MLLM unlearning , enabling test-time behavior control without parameter updates.
@@ 2.1 Notation and Problem Setup
where $\lambda_{f},\lambda_{r}$ are trade-off parameters. The retain loss $\mathcal{L}_{r}$ is commonly instantiated as an autoregressive negative log-likelihood (NLL) or a KL-divergence constraint term (Liu et al., 2025c), ensuring that the model preserves performance on $\mathcal{D}_{r}$. The forget loss $\mathcal{L}_{f}$ serves as the unlearning objective, typically implemented through gradient ascent (Thudi et al., 2022) or preference-based optimization (Zhang et al., 2024), encouraging the model to deviate from its original predictions.
@@ 2.2 Behavioral Control through Activation Steering
$\mathbf{d}_{\text{erase}}=\frac{1}{|\mathcal{D}^{+}|}\sum_{(\mathcal{I},\mathcal{Q})\in\mathcal{D}^{+}}\mathbf{h}^{\ell}(\mathcal{I},\mathcal{Q})-\frac{1}{|\mathcal{D}^{-}|}\sum_{(\mathcal{I},\mathcal{Q})\in\mathcal{D}^{-}}\mathbf{h}^{\ell}(\mathcal{I},\mathcal{Q}),$ (3)
where $\mathcal{D}^{+}$ denotes the set of knowledge-recall samples and $\mathcal{D}^{-}$ the corresponding knowledge-erasure samples, which will be discussed in Section 3.1. The resulting direction is subsequently added to the hidden states to steer the representation, formulated as:
@@ 3.1 Multimodal Erasure Direction Construction
$\mathcal{I}_{i}^{(k+1)^{\prime}}\,=\Pi_{\mathcal{I}_{i}+\mathcal{B}}\,\!\Big(\mathcal{I}_{i}^{(k)^{\prime}}+\alpha\cdot\mathrm{sign}\,\Big(\nabla_{\mathcal{I}}\log P_{\theta}\big(y\mid\mathcal{I}^{(k)^{\prime}}_{i},\mathcal{Q}_{i}\big)\,\Big)\Big),$ (6)
@@
$\mathbf{d}_{\text{erase}}^{\ell}=\frac{1}{|\mathcal{D}^{+}|}\sum_{(\mathcal{I},\mathcal{Q})\in\mathcal{D}^{+}}\mathbf{h}(\mathcal{I},\mathcal{Q})-\frac{1}{|\mathcal{D}^{-}|}\sum_{(\mathcal{I^{\prime}},\mathcal{Q^{\prime}})\in\mathcal{D}^{-}}\mathbf{h}(\mathcal{I^{\prime}},\mathcal{Q^{\prime}}).$ (7)
@@ 3.2 Input-aware Steering with Erasure Directions
After getting the multimodal erasure direction, it is essential to determine when to apply it, so that it would not affect the model performance on the retain set. Current steering methods often indiscriminately apply the steering vector to all prompts (Rimsky et al., 2024), a strategy that inevitably degrades overall model performance. Regarding unlearning, this degradation manifests as corrupted responses to samples in the retain set, leading to over-forgetting problem (Xu et al., 2025). An desirable unlearning steering mechanism should be input-aware: for samples in the forget set, their activations should be steered toward the knowledge-erasure behavior, while for samples in the retain set, their activations should remain as unchanged as possible. To achieve this, the steering task can be formulated as an input-aware task by constructing a direction-determining function $f(\mathbf{h}(\mathcal{I},\mathcal{Q}))$, which conditions on the query’s activation and produces the steering direction applied at inference. Then the steering process can be written as:
@@ 4.3 Ablation Study of MLLMEraser (RQ2)
To assess the effectiveness of our proposed multimodal erasure direction construction and input-aware steering mechanism, we introduce two variants: Text-only erasure direction, which derives $\mathbf{d}_{\text{erase}}$ solely from refusal/jailbreak text pairs without incorporating visual information, and Input-unaware steering, which applies the erasure direction uniformly to all inputs without selective control. The results are summarized in Table 2, and we further visualize how the activation distributions of the forget and retain sets shift before and after unlearning in Figure 4. We can observe that:
@@ 4.4 Results Analysis for Unlearning Efficiency (RQ3)
We further evaluate the efficiency of different unlearning methods in terms of both training and inference time, as shown in Figure 5. Details on GPU memory usage can be found in Appendix E. Here, training time refers to the total time required to obtain an unlearned model, while inference time indicates the total time spent processing $10$ randomly sampled inputs. We can find that:
@@
Overall, MLLMEraser provides a lightweight framework for test-time unlearning in MLLMs, avoiding parameter updates while introducing only negligible test-time overhead.
@@ C.1 Activation Steering
In multimodal models, Gan et al. (2025) transfer language-derived vectors to enhance visual reasoning, VTI (Liu et al., 2025b) and L2S (Parekh et al., 2025) mitigate hallucinations through input-dependent interventions, and AutoSteer (Wu et al., 2025) automates safe steering. Additional works explore modality preference steering (Zhang et al., 2025b) and analyze how finetuning reshapes steerable representations (Khayatan et al., 2025).
@@ C.2 LLM Unlearning
The problem of unlearning in large language models (LLMs) has attracted increasing attention due to growing concerns over privacy leakage, copyright infringement, and safety risks. Early approaches primarily relied on gradient-ascent (Thudi et al., 2022) fine-tuning, which attempts to maximize the loss on samples to be forgotten so as to erase their influence (Liu et al., 2024b; Maini et al., 2024). While conceptually simple, these methods were quickly shown to be unstable, often leading to catastrophic degradation of model utility across retain data. To overcome these limitations, subsequent research proposed more principled optimization frameworks, such as preference-based unlearning (Zhang et al., 2024), weight-saliency–driven parameter editing (Fan et al., 2024), and pruning-oriented removal of knowledge (Zhang et al., 2025a). Parameter-efficient strategies (Ding et al., 2025; Liu et al., 2024c) further reduced the overhead compared with full-model finetuning, while unified gradient-based formulations (Huang et al., 2024) and knowledge-gap alignment methods (Liu et al., 2024b; Wang et al., 2023) aimed to improve stability and generalization. Beyond optimization-centric methods, continual private unlearning settings (Liu et al., 2022a) extend the scope to dynamic data distributions, while neuron-level editing (Wu et al., 2023) and copyright-specific takedown mechanisms (Dou et al., 2025) address more fine-grained or domain-driven requirements. In parallel, lightweight alternatives such as guardrail prompting (Thaker et al., 2024), in-context unlearning (Pawelczyk et al., 2024), and task-vector editing (Ilharco et al., 2023) illustrate that test-time or post-hoc interventions can also provide partial forgetting.
@@ C.3 MLLM Unlearning
Compared with LLMs, research on unlearning in multimodal large language models (MLLMs) is still nascent. A first line of work investigates vision–language models such as CLIP (Radford et al., 2021). CLIPErase (Yang et al., 2025) develops forgetting–retention–consistency objectives to selectively erase visual–textual associations while preserving unrelated semantics. In the generative domain, Erasing Concepts from Diffusion Models (Fuchi & Takagi, 2024) demonstrates that fine-tuning diffusion weights can remove high-level concepts (e.g., nudity, artistic styles) while maintaining unrelated generative capabilities.
For fully multi-modal architectures, existing approaches can be broadly grouped into two categories. (i) Direct migrations of LLM unlearning methods, where objectives such as gradient ascent (Thudi et al., 2022), NPO (Zhang et al., 2024), or KL-based formulations are adapted into multi-modal finetuning pipelines (Huo et al., 2025; Liu et al., 2025d). Cross-Modal Safety Alignment (Chakraborty et al., 2024) further shows that even textual unlearning alone, applied at the LLM backbone, can effectively transfer to vision–language models and substantially reduce multi-modal jailbreak success rates, offering a cost-efficient alternative to full multimodal finetuning. (ii) Selective or architecture-aware updates, which target specific parameters to mitigate side effects. Single Image Unlearning (SIU) (Li et al., 2024) addresses the challenge of forgetting visual concepts with limited data by introducing a Dual Masked KL-divergence (DMK) Loss, which applies token-level and vocabulary-level masking to decouple factual knowledge from visual recognition and preserve non-target knowledge. MMUnlearner (Huo et al., 2025) advances this direction by leveraging weight saliency and geometric constraints to erase visual traces while retaining textual information, while MANU (Liu et al., 2025d) introduces modality-aware neuron pruning to balance forgetting across modalities. EFUF (Xing et al., 2024) further applies fine-grained gradient-ascent unlearning to reduce multimodal hallucinations by selectively editing spurious visual features.
@@ D.3 Baseline Methods
NPO (Zhang et al., 2024) formulates unlearning as a variant of preference optimization without positive examples. Forget set samples are treated as dispreferred responses, and the loss penalizes their probability relative to a reference model trained only on $\mathcal{D}_{r}$. The objective is:
@@
D.4 Implementation Details.
@@
The vanilla and baseline models are implemented following the configurations reported in their original papers (Liu et al., 2025c; Huo et al., 2025), ensuring consistency with prior unlearning studies. For both LLaVA-1.5 and Qwen-2.5-VL models, we adopt LoRA during fine-tuning to reduce memory usage. For our proposed method, the steering strength $\lambda$ is set to $0.3$ and the regularization parameter $\gamma=1.0$ on LLaVA-1.5-7B, while on Qwen-2.5-VL-7B we use $\lambda=0.25$ and $\gamma=0.1$. For the construction of harmful textual data, we follow the setting in (Zhao et al., 2025) to construct the textual erasure direction. For adversarial visual samples, the clean images are sampled from ImageNet (Deng et al., 2009) and perturbation radius is set to $\epsilon=16/255$.
Appendix E Discussion About the Efficiency of MLLMEraser

### v3

@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023; Wang et al., 2024; Zhu et al., 2024; Yang et al., 2023; Anil et al., 2023) have demonstrated remarkable capabilities in tasks such as visual question answering (Hu et al., 2024; Kuang et al., 2025), image–text generation (Wu et al., 2024b; Lan et al., 2025), and embodied AI applications (Wu et al., 2024a; Cheng et al., 2025). However, their large-scale deployment raises concerns about memorizing problematic information once learned, particularly in privacy-sensitive (Huo et al., 2025) or safety-critical (Liu et al., 2024c) applications, highlighting the need for reliable unlearning mechanisms to ensure trustworthy MLLM systems (Li et al., 2024; Liu et al., 2025c; Chen et al., 2025). MLLM unlearning aims to selectively erase designated information across modalities while preserving general utility, thereby supporting privacy protection, mitigating misuse, and maintaining reliability (Dontsov et al., 2025). Existing work mainly adapts training–based strategies from LLM unlearning, employing gradient ascent, preference optimization (Zhang et al., 2024a), or performing targeted parameter updates (Liu et al., 2025d; Huo et al., 2025). While effective, these training-based methods introduce substantial computational costs, inference latency, and risks of corrupting retained knowledge (Ding et al., 2025). This motivates test-time MLLM unlearning (See Figure 1), a paradigm that prevents the generation of designated information at inference without modifying model parameters, which offers an immediate, lightweight, and reversible solution.
@@ 2 Preliminary
This section begins by introducing the notation and formalizing the problem of test-time MLLM unlearning in Section 2.1. Section 2.2 outlines how activation steering offers a feasible mechanism for test-time MLLM unlearning, enabling test-time behavior control.
@@ 2.1 Notation and Problem Setup
[…]
where $\lambda_{f},\lambda_{r}$ are trade-off parameters. The retain loss $\mathcal{L}_{r}$ is commonly instantiated as an autoregressive negative log-likelihood (NLL) or a KL-divergence constraint term (Liu et al., 2025c), ensuring that the model preserves performance on $\mathcal{D}_{r}$. The forget loss $\mathcal{L}_{f}$ serves as the unlearning objective, typically implemented through gradient ascent (Thudi et al., 2022) or preference-based optimization (Zhang et al., 2024a), encouraging the model to deviate from its original predictions.
@@ 2.2 Behavioral Control through Activation Steering
$\mathbf{d}_{\text{erase}}=\frac{1}{|\mathcal{D}^{+}|}\!\!\!\!\!\!\sum_{\hskip 8.19447pt(\mathcal{I},\mathcal{Q})\in\mathcal{D}^{+}}\!\!\!\!\!\!\mathbf{h}^{\ell}(\mathcal{I},\mathcal{Q})-\frac{1}{|\mathcal{D}^{-}|}\!\!\!\!\!\!\sum_{\hskip 8.19447pt(\mathcal{I},\mathcal{Q})\in\mathcal{D}^{-}}\!\!\!\!\!\!\mathbf{h}^{\ell}(\mathcal{I},\mathcal{Q}),$ (3)
where $\mathcal{D}^{+}$ denotes the set of knowledge-erasure samples and $\mathcal{D}^{-}$ the corresponding knowledge-recall samples, which will be discussed in Section 3.1. The resulting direction is subsequently added to the hidden states to steer the representation, formulated as:
@@ 3.1 Multimodal Erasure Direction Construction
$\mathcal{I}_{i}^{(k+1)^{\prime}}\!=\!\prod_{\mathcal{I}_{i}\!+\mathcal{B}}\!\Big(\mathcal{I}_{i}^{(k)^{\prime}}\!+\alpha\cdot\mathrm{sign}\Big(\nabla_{\mathcal{I}}\log P_{\theta}\big(y\mid\mathcal{I}^{(k)^{\prime}}_{i}\!,\mathcal{Q}_{i}\big)\Big)\Big),$ (6)
@@
$\mathbf{d}_{\text{erase}}^{\ell}=\frac{1}{|\mathcal{D}^{+}|}\!\!\!\!\!\!\sum_{\hskip 8.19447pt(\mathcal{I},\mathcal{Q})\in\mathcal{D}^{+}}\!\!\!\!\!\!\mathbf{h}(\mathcal{I},\mathcal{Q})-\frac{1}{|\mathcal{D}^{-}|}\!\!\!\!\!\!\sum_{\hskip 8.19447pt(\mathcal{I^{\prime}},\mathcal{Q^{\prime}})\in\mathcal{D}^{-}}\!\!\!\!\!\!\!\!\!\mathbf{h}(\mathcal{I^{\prime}},\mathcal{Q^{\prime}}).$ (7)
@@ 3.2 Input-aware Steering with Erasure Directions
After getting the multimodal erasure direction, it is essential to determine when to apply it, so that it would not affect the model performance on the retain set. Current steering methods often indiscriminately apply the steering vector to all prompts (Rimsky et al., 2024), a strategy that inevitably degrades overall model performance. Regarding unlearning, this degradation manifests as corrupted responses to samples in the retain set, leading to over-forgetting problem (Xu et al., 2025). An desirable unlearning steering mechanism should be input-aware: for samples in the forget set, their activations should be steered toward the knowledge-erasure behavior, while for samples in the retain set, their activations should remain as unchanged as possible.
To achieve this, the steering task can be formulated as an input-aware task by constructing a direction-determining function $f(\mathbf{h}(\mathcal{I},\mathcal{Q}))$, which conditions on the query’s activation and produces the steering direction applied at inference. Then the steering process can be written as:
@@ 4.3 Ablation Study of MLLMEraser (RQ2)
To assess the effectiveness of our proposed multimodal erasure direction construction and input-aware steering mechanism, we introduce two variants: Text-only erasure direction, which derives $\mathbf{d}_{\text{erase}}$ solely from refusal/jailbreak text pairs without incorporating visual information, and Input-unaware steering, which applies the erasure direction uniformly to all inputs without selective control. The results are presented in Table 2, and we further visualize how the activation distributions of the forget and retain sets shift before and after unlearning in Figure 8. We can observe that:
@@ 4.4 Results Analysis for Unlearning Efficiency (RQ3)
We further evaluate the efficiency of different unlearning methods in terms of both training and inference time, as shown in Figure 5. Details on GPU memory usage can be found in Appendix F. Here, training time refers to the total time required to obtain an unlearned model, while inference time indicates the total time spent processing $10$ randomly sampled inputs. We can find that:
@@
Overall, MLLMEraser provides a lightweight framework for MLLM test-time unlearning, avoiding parameter updates while introducing only negligible test-time overhead.
@@ C.1 Activation Steering
In multimodal models, (Gan et al., 2025) transfer language-derived vectors to enhance visual reasoning, VTI (Liu et al., 2025b) and L2S (Parekh et al., 2025) mitigate hallucinations through input-dependent interventions, and AutoSteer (Wu et al., 2025) automates safe steering. Additional works explore modality preference steering (Zhang et al., 2025c) and analyze how finetuning reshapes steerable representations (Khayatan et al., 2025).
@@ C.2 LLM Unlearning
The problem of unlearning in large language models (LLMs) has attracted increasing attention due to growing concerns over privacy leakage, copyright infringement, and safety risks. Early approaches primarily relied on gradient-ascent (Thudi et al., 2022) fine-tuning, which attempts to maximize the loss on samples to be forgotten so as to erase their influence (Liu et al., 2024b; Maini et al., 2024). While conceptually simple, these methods were quickly shown to be unstable, often leading to catastrophic degradation of model utility across retain data. To overcome these limitations, subsequent research proposed more principled optimization frameworks, such as preference-based unlearning (Zhang et al., 2024a), weight-saliency–driven parameter editing (Fan et al., 2024), and pruning-oriented removal of knowledge (Zhang et al., 2025a). Parameter-efficient strategies (Ding et al., 2025; Liu et al., 2024c) further reduced the overhead compared with full-model finetuning, while unified gradient-based formulations (Huang et al., 2024) and knowledge-gap alignment methods (Liu et al., 2024b; Wang et al., 2023) aimed to improve stability and generalization. Beyond optimization-centric methods, continual private unlearning settings (Liu et al., 2022a) extend the scope to dynamic data distributions, while neuron-level editing (Wu et al., 2023) and copyright-specific takedown mechanisms (Dou et al., 2025) address more fine-grained or domain-driven requirements. In parallel, lightweight alternatives such as guardrail prompting (Thaker et al., 2024), in-context unlearning (Pawelczyk et al., 2024), and task-vector editing (Ilharco et al., 2023) illustrate that test-time or post-hoc interventions can also provide partial forgetting.
@@ C.3 MLLM Unlearning
Compared with LLMs, research on unlearning in multimodal large language models (MLLMs) is still nascent. A first line of work investigates vision–language models such as CLIP (Radford et al., 2021). CLIPErase (Yang et al., 2025) develops forgetting–retention–consistency objectives to selectively erase visual–textual associations while preserving unrelated semantics. In the generative domain, Erasing Concepts from Diffusion Models (Fuchi and Takagi, 2024) demonstrates that fine-tuning diffusion weights can remove high-level concepts (e.g., nudity, artistic styles) while maintaining unrelated generative capabilities.
For fully multi-modal architectures, existing approaches can be broadly grouped into two categories. (i) Direct migrations of LLM unlearning methods, where objectives such as gradient ascent (Thudi et al., 2022), NPO (Zhang et al., 2024a), or KL-based formulations are adapted into multi-modal finetuning pipelines (Huo et al., 2025; Liu et al., 2025d). Cross-Modal Safety Alignment (Chakraborty et al., 2024) further shows that even textual unlearning alone, applied at the LLM backbone, can effectively transfer to vision–language models and substantially reduce multi-modal jailbreak success rates, offering a cost-efficient alternative to full multimodal finetuning. (ii) Selective or architecture-aware updates, which target specific parameters to mitigate side effects. Single Image Unlearning (SIU) (Li et al., 2024) addresses the challenge of forgetting visual concepts with limited data by introducing a Dual Masked KL-divergence (DMK) Loss, which applies token-level and vocabulary-level masking to decouple factual knowledge from visual recognition and preserve non-target knowledge. MMUnlearner (Huo et al., 2025) advances this direction by leveraging weight saliency and geometric constraints to erase visual traces while retaining textual information, while MANU (Liu et al., 2025d) introduces modality-aware neuron pruning to balance forgetting across modalities. EFUF (Xing et al., 2024) further applies fine-grained gradient-ascent unlearning to reduce multimodal hallucinations by selectively editing spurious visual features.
@@ D.3 Baseline Methods
NPO (Zhang et al., 2024a) formulates unlearning as a variant of preference optimization without positive examples. Forget set samples are treated as dispreferred responses, and the loss penalizes their probability relative to a reference model trained only on $\mathcal{D}_{r}$. The objective is:
@@
MANU.
MANU (Liu et al., 2025d) performs unlearning by selectively pruning neurons that contribute more to the forget set than to the retain set. The method first computes modality-aware neuron importance using activation statistics across multimodal and textual inputs, and then assigns each neuron a pruning score reflecting its relative contribution to forgotten knowledge. Neurons with the highest scores are pruned, enabling targeted removal of undesired multimodal behavior while minimizing disruption to retained capabilities.
D.4 Implementation Details
@@
The vanilla and baseline models are implemented following the configurations reported in their original papers (Liu et al., 2025c; Huo et al., 2025), ensuring consistency with prior unlearning studies. For both LLaVA-1.5 and Qwen-2.5-VL models, we adopt LoRA during fine-tuning to reduce memory usage. For our proposed method, the steering strength $\lambda$ is set to $0.3$ and the regularization parameter $\gamma=1.0$ on LLaVA-1.5-7B, while on Qwen-2.5-VL-7B we use $\lambda=0.25$ and $\gamma=0.1$. All experiments are conducted on NVIDIA A800 GPUs (80 GB). For the construction of harmful textual data, we follow the setting in (Zhao et al., 2025) to construct the textual erasure direction. For adversarial visual samples, the clean images are sampled from ImageNet (Deng et al., 2009) and perturbation radius is set to $\epsilon=16/255$.
Appendix E Hyperparameter Analysis
In this section, we provide a comprehensive analysis of the key hyperparameters in MLLMEraser, namely the regularization parameter $\gamma$, the steering strength $\lambda$, and the perturbation budget $\epsilon$.
These hyperparameters govern different aspects of the test-time unlearning process: $\gamma$ acts as a regularization term, $\lambda$ determines the magnitude of the steering intervention, and $\epsilon$ specifies the radius for constructing the erasure direction.
E.1 Regularization Parameter $\gamma$
The corresponding results are presented in Figure 6. When $\gamma$ is small, the regularization term plays only a mild role. It constrains the erasure direction just enough to prevent overfitting, but not strongly enough to interfere with the forgetting objective. As a result, the method can still focus on the distinctive activation differences between the forget and retain samples, leading to stable and robust performance across small values of $\gamma$.
However, when $\gamma$ becomes too large (e.g., $\gamma=10$), the regularization begins to dominate the optimization. In this case, the method is overly restricted and becomes reluctant to modify the activations associated with the forget set. This suppresses the useful forgetting signal extracted from the contrastive pairs and forces the learned direction to remain too close to the retain set’s behavior. Consequently, the erasure effect becomes significantly weaker, leading to noticeably worse forgetting performance.
(a) Results on the classification task.
(b) Results on the generation task.
(c) Results on the cloze task.
Figure 6: Sensitivity analysis of the regularization parameter $\gamma$ on LLaVA-1.5-7B under the 10% forgetting setting. 6(a) reports results on the classification task, 6(b) shows results on the generation task, and 6(c) presents results on the cloze task.
E.2 Steering Strength $\lambda$
The corresponding results are presented in Figure 7. We further tune $\lambda$ within $[0.1,0.15,0.20,0.25,0.30,0.35]$. Increasing $\lambda$ consistently strengthens the erasure effect, while the model utility remains largely unaffected. This behavior is expected: a larger $\lambda$ amplifies the steering vector, pushing forget-set activations more aggressively toward the erasure direction. As long as $\lambda$ remains within a moderate range, the retain-set activations stay mostly within the original subspace, and thus their semantics are preserved. Only when $\lambda$ becomes excessively large do we observe slight utility degradation, suggesting that over-steering begins to distort general representations. Overall, these findings highlight the advantage of the null-space projection constraint, which provides a wide operational range where stronger forgetting does not compromise model utility.
(a) Results on the classification task.
(b) Results on the generation task.
(c) Results on the cloze task.
Figure 7: Sensitivity analysis of the steering strength $\lambda$ on LLaVA-1.5-7B under the 10% forgetting setting. 7(a) reports results on the classification task, 7(b) shows results on the generation task, and 7(c) presents results on the cloze task.
For our proposed method, the steering strength $\lambda$ is set to $0.3$ and the regularization parameter $\gamma=1.0$ on LLaVA-1.5-7B, while on Qwen-2.5-VL-7B we use $\lambda=0.25$ and $\gamma=0.1$.
E.3 Perturbation Budget $\epsilon$
The corresponding results are presented in Table 3. We observe that when $\epsilon$ is small, the method achieves both strong forgetting performance and high utility preservation. As $\epsilon$ increases, the steering vector gains stronger knowledge-erasure capability, resulting in stronger forgetting. However, when $\epsilon$ becomes too large, the erasure directions begin to encode spurious noise rather than meaningful semantic differences, leading to degraded performance in both forgetting quality and model utility. We set the perturbation radius to $\epsilon=\frac{16}{255}$, which provides a sufficiently expressive search region for extracting an effective erasure direction while avoiding overly noisy gradients.
Table 3: Unlearning performance on MLLMU-Bench (10% Forget) under different perturbation budgets $\epsilon$. Results are evaluated on the forget set (Fgt), test set (Test), retain set (Ret), and celebrity set (Cele). $\downarrow$ indicates lower is better, and $\uparrow$ indicates higher is better.
LLaVA-1.5-7B (10% Forget)-VQA
$\epsilon=\frac{16}{255}$ 2.00 27.20 45.08 46.48 0.176 0.202 0.594 0.314 5.00 6.00 54.33 7.84
$\epsilon=\frac{32}{255}$ 0.00 14.00 42.94 44.26 0.040 0.133 0.584 0.308 2.00 6.00 52.56 6.86
$\epsilon=\frac{64}{255}$ 2.00 20.40 44.59 46.21 0.355 0.228 0.593 0.314 14.00 12.00 53.56 7.19
unstrained 6.00 24.40 44.32 46.21 0.355 0.228 0.593 0.314 5.00 6.00 54.33 7.84
Appendix F Discussion About the Efficiency of MLLMEraser